# Mixmaster chaos in an AdS black hole interior

### v1

@@ Abstract
We derive gravitational backgrounds that are asymptotically Anti-de Sitter, have a regular black hole horizon and which deep in the interior exhibit mixmaster chaotic dynamics. The solutions are obtained by coupling gravity with a negative cosmological constant to three massive vector fields, within an ansatz that reduces to ordinary differential equations. At late interior times the equations are identical to those analyzed in depth by Misner and by Belinskii-Khalatnikov-Lifshitz fifty years ago. We review and extend known classical and semiclassical results on the interior chaos, formulated as both a dynamical system of ‘Kasner eras’ and as a hyperbolic billiards problem. The volume of the universe collapses doubly-exponentially over each Kasner era. A remarkable feature is the emergence of a conserved energy, and hence a ‘time-independent’ Hamiltonian, at asymptotically late interior times. A quantisation of this Hamiltonian exhibits arithmetic chaos associated with the principal congruence subgroup $\Gamma(2)$ of the modular group. We compute a large number of eigenvalues numerically to obtain the spectral form factor. While the spectral statistics is anomalous for a chaotic system, the eigenfunctions themselves display random matrix behaviour.
@@ 1 Introduction
Kasner singularities are characterised by their scaling exponents, and hence one might hope to identify these exponents in the dual field theory. Complicating this picture, subsequent work rediscovered, in a holographic context, the well-known phenomenon of Kasner bounces, in which the Kasner exponents abruptly change [18, 19, 20, 21]. See also [22]. This leads to the notion of a Kasner epoch, of finite duration. Several holographic works found scenarios in which the sequence of epochs connected by bounces continues indefinitely, such that a final asymptotic Kasner regime is never reached [23, 24, 25, 26, 27]. These studies were able numerically to evolve the dynamics through a number of Kasner epochs. However, as was understood in the original work from the Landau Institute [1, 2], reviewed in detail in §3 and §4 below, the chaotic nature of the dynamics is revealed in the very long-term evolution of so-called Kasner eras. A Kasner era is a potentially large number of Kasner epochs that involves bounces between only two neighbouring ‘walls’. Numerical evolution through multiple Kasner eras is not straightforward (although see [28]).
@@
The same mathematical structure, harmonic analysis on a modular domain, has recently been used to study two-dimensional CFT partition functions [43]. Arithmetic quantum chaos may provide a useful perspective on CFT quantum chaos [44, 45]. Furthermore, the modular subgroup $\Gamma(2)$ has previously appeared in the context of the CFT bootstrap [46]. It would be very interesting, of course, to establish any kind of connection between the near-singularity gravitational dynamics and scrambling in the dual CFT. We should emphasise, however, to avoid potential confusion, that the chaos that emerges towards gravitational singularities is distinct from the chaos associated to the infinite blueshift that can be experienced by an observer close to a horizon, which has been studied in depth in recent years following [47].
@@ 2 Setup with massive vector fields
The cosmological constant ensures the AdS asymptotics, with unit AdS radius. Near the singularity the mass terms and the cosmological constant will become subleading. In this far interior regime, without the mass terms, the field strengths are fixed by a Gauss law while producing ‘electric walls’ that bound the evolution of the metric components [48, 49]. We will need three Maxwell fields to fully bound the metric components and ensure chaotic dynamics towards the singularity. While the mass terms drop out towards the singularity, they are necessary for this theory to have a planar black hole exterior solution with all three vector fields nonzero, within the Ansatz that we now introduce. The masses will also ensure the absence of Cauchy horizons in the interior. We will choose the masses of the vector fields to be such that the dual vector operators are relevant deformations of the CFT. Sourcing this operator therefore preserves the AdS asymptotics.
@@
Here again the $\phi_{i}$ are functions of $z$. Substituting the Ansatz into the equations of motion,
@@
For the purposes of giving numerical illustrations of our setup we will set the three masses equal with the choice $\mu_{i}^{2}=-\frac{3}{16}$. With this choice of negative mass squared, which is stable in AdS spacetime, the asymptotic behavior of the fields as $z\to 0$ at the AdS boundary is
@@
The fact that the leading source term $\phi_{i}^{(0)}$ goes to zero as $z\to 0$ ensures the survival of the AdS asymptotics.
@@ 3.3 The end of a Kasner epoch
[…]
@@
These constraints follow from (12), (22) and (23). The following Fig. 4 shows three examples of the bounded evolution of the universe.
Figure 4: The late time interior dynamics of $g(\rho)$ and $h(\rho)$ is confined to the triangle (24). The plots show three different illustrative trajectories, generated numerically using the full equations of motion (4) and (5). The straight lines correspond to Kasner epochs. An era is given by a sequence of epochs bouncing between two given walls. The leftmost plot is zoomed in to a tip of the triangle, showing an era of many epochs. Left: $\phi_{t}^{(0)}\approx 1.49011\,T^{5/4}$, $\phi_{x}^{(0)}=3\phi_{t}^{(0)}$ and $\phi_{y}^{(0)}=4\phi_{t}^{(0)}$. Middle: $\phi_{t}^{(0)}\approx 2.88867\,T^{5/4}$, $\phi_{x}^{(0)}=2\phi^{(0)}_{t}$ and $\phi_{y}^{(0)}=4\phi^{(0)}_{t}$. Right: $\phi_{t}^{(0)}\approx 5.50918\,T^{5/4}$, $\phi_{x}^{(0)}=2\phi_{t}^{(0)}$ and $\phi_{y}^{(0)}=3\phi_{t}^{(0)}$.
It is clear from the figure that each Kasner epoch ends in a ‘bounce’ to a subsequent Kasner epoch. We will characterise such bounces in detail in the following sections.
While the potential (11) is not identical to the original Bianchi IX mixmaster potential from [1, 2, 3, 4], the walls in (24) are the same as the dominant walls in that case. That is to say, the additional walls in the Bianchi IX dynamics are outside of the triangle (24) and are therefore irrelevant for the asymptotic interior dynamics. This is why we will recover the known recursion relations between Kasner epochs.
@@ 3.4 Reflection rule
[…]
@@ 4.1 Era dynamics
[…]
@@ 5.1 The Wheeler-DeWitt equation
Here we have again considered late interior times so that the mass and cosmological constant terms in the action can be dropped.
@@
As commented in footnote 2 above, the minisuperspace WDW equation is only trustworthy to leading order in the semiclassical limit. Beyond that limit fluctuations of inhomogeneous modes contribute to the wavefunction and ordering ambiguities become relevant. Restricting, then, to highly excited states one has from (77) that
@@ 5.3 Congruence subgroups and automorphic forms
[…]
@@ 6.1 Numerical method to determine large collections of eigenvalues
[…]
@@
As previously, the matrix $\mathbf{B}({\varepsilon}_{n})$, with components $B_{lm}({\varepsilon}_{n})$, is ill-conditioned. Finding suitable pre-conditioners might appear challenging, as the boundary condition (108) involves multiple Bessel functions that need to balance against each other. However, after some trial and error we have found that it is sufficient to define a matrix $\mathbf{B}^{\prime}({\varepsilon}_{n})$ just as in (106b), but with $m$ in (106b) replaced by $(m-1/2)$. For the Neumann case, computing the marix $\mathbf{B}^{\prime}({\varepsilon}_{n})$ is more daunting, and we have only managed to accurately determine the first 1200 eigenvalues. Again, below, we will provide tests of our data suggesting that we have captured all the Neumann eigenvalues in the range $0<\varepsilon_{n}\lesssim 170.31(7)$.
@@ 6.2 Numerical scheme for finding $\varepsilon_{n}$ and few thousand $c^{n}_{m}$
With a given $\varepsilon_{n}$ known to some precision, we can change gears and use a different numerical method to determine $\varepsilon_{n}$ and the corresponding $c^{n}_{m}$ more accurately. In this method one solves for the eigenpair $\{\Psi_{n},\varepsilon_{n}\}$ directly on a discretised grid. First, we change to a new set of coordinates wherein $(x,y)$ are mapped to coodinates $(X,Y)$,
[…]
@@
[…]
@@ 6.3 Validating the numerical data
where the subscript ${}_{D}$ is for ‘Dirichlet’. The leading term in the expansion is determined by the area of the billiard and is independent of the boundary conditions or shape of the billiard. [^7] Locally one may consider the Laplace equation to be in flat space with eigenstates $e^{ik\cdot x}$ and eigenvalues $\mu=k^{2}$. Given a region of area $A$ one can, by the usual counting of plane wave states, fit $N=A\int d^{2}k/(2\pi)^{2}$ states into the region. Setting $d^{2}k=2\pi kdk$ and integrating up to $k=\sqrt{\mu}$ gives $N(\mu)=\frac{A}{4\pi}\mu$. The area of the small triangle defined by the interior of (96) is simply $\pi/6$.
@@
where the subscript ${}_{N}$ is to ‘Neumann’ and $\{A_{0},B_{0},C_{0}\}$ are to be fit. The first term in (114) is the same as for the Dirichlet sector. A simple $\chi^{2}$ fit yields
[…]
@@ 7.1 Density of states
where $A=\pi/6$ is the area of the small triangle defined in (96). We recalled the origin of this leading term in footnote 7 above. Although we shall work at fixed energy throughout, we can make a brief comment about the canonical ensemble. The density of states (117) corresponds to the entropy
@@ 7.4 Hecke operators
The matrices in (132) do not belong to $\Gamma(2)$, and hence it is not immediate that the action (131) maps the space of $\Gamma(2)$ automorphic forms to itself. We will now establish that fact.
@@
[…]
@@ 7.5 Waveform coefficients
[…]

### v4

@@ Abstract
We derive gravitational backgrounds that are asymptotically Anti-de Sitter, have a regular black hole horizon and which deep in the interior exhibit mixmaster chaotic dynamics. The solutions are obtained by coupling gravity with a negative cosmological constant to three massive vector fields, within an Ansatz that reduces to ordinary differential equations. At late interior times the equations are identical to those analysed in depth by Misner and by Belinskii-Khalatnikov-Lifshitz fifty years ago. We review and extend known classical and semiclassical results on the interior chaos, formulated as both a dynamical system of ‘Kasner eras’ and as a hyperbolic billiards problem. The volume of the universe collapses doubly-exponentially over each Kasner era. A remarkable feature is the emergence of a conserved energy, and hence a ‘time-independent’ Hamiltonian, at asymptotically late interior times. A quantisation of this Hamiltonian exhibits arithmetic chaos associated with the principal congruence subgroup $\Gamma(2)$ of the modular group. We compute a large number of eigenvalues numerically to obtain the spectral form factor. While the spectral statistics is anomalous for a chaotic system, the eigenfunctions themselves display random matrix behaviour.
@@ 1 Introduction
Kasner singularities are characterised by their scaling exponents, and hence one might hope to identify these exponents in the dual field theory. Complicating this picture, subsequent work rediscovered, in a holographic context, the well-known phenomenon of Kasner bounces, in which the Kasner exponents abruptly change [18, 19, 20, 21]. See also [22]. This leads to the notion of a Kasner epoch, of finite duration. Several holographic works found scenarios in which the sequence of epochs connected by bounces continues indefinitely, such that a final asymptotic Kasner regime is never reached [23, 24, 25, 26, 27, 28]. These studies were able numerically to evolve the dynamics through a number of Kasner epochs. However, as was understood in the original work from the Landau Institute [1, 2], reviewed in detail in §3 and §4 below, the chaotic nature of the dynamics is revealed in the very long-term evolution of so-called Kasner eras. A Kasner era is a potentially large number of Kasner epochs that involves bounces between only two neighbouring ‘walls’. Numerical evolution through multiple Kasner eras is not straightforward (although see [29]).
@@
The same mathematical structure, harmonic analysis on a modular domain, has recently been used to study two-dimensional CFT partition functions [44]. Arithmetic quantum chaos may provide a useful perspective on CFT quantum chaos [45, 46]. Furthermore, the modular subgroup $\Gamma(2)$ has previously appeared in the context of the CFT bootstrap [47]. It would be very interesting, of course, to establish any kind of connection between the near-singularity gravitational dynamics and scrambling in the dual CFT. We should emphasise, however, to avoid potential confusion, that the chaos that emerges towards gravitational singularities is distinct from the chaos associated to the infinite blueshift that can be experienced by an observer close to a horizon, which has been studied in depth in recent years following [48]. However, similar chaotic can also arise towards a timelike singularity, and this may potentially be more accessible to the boundary theory [49].
@@ 2 Setup with massive vector fields
The cosmological constant ensures the AdS asymptotics, with unit AdS radius. Near the singularity the mass terms and the cosmological constant will become subleading. In this far interior regime, without the mass terms, the field strengths are fixed by a Gauss law while producing ‘electric walls’ that bound the evolution of the metric components [50, 51]. We will need three Maxwell fields to fully bound the metric components and ensure chaotic dynamics towards the singularity. Although will not pursue a string-theoretic realisation of (1), we may note that multiple Maxwell fields arise, for example, in gauged supergravity theories [52]. While the mass terms drop out towards the singularity, they are necessary for this theory to have a planar black hole exterior solution with all three vector fields nonzero, within the Ansatz that we now introduce. The masses will also ensure the absence of Cauchy horizons in the interior. We will choose the masses of the vector fields to be such that the dual vector operators are relevant deformations of the CFT. Sourcing this operator therefore preserves the AdS asymptotics.
@@
Here again the $\phi_{i}$ are functions of $z$. It is important that (3) does not include a longitudinal component, i.e. $\vec{\nabla}\cdot\vec{A}_{i}=0$ on the constant $z$ slices, as a longitudinal component is not bounded by electric walls [24]. Substituting the Ansatz into the equations of motion,
@@
For the purposes of giving numerical illustrations of our setup we will set the three masses equal with the choice $\mu_{i}^{2}=-\frac{3}{16}$. With this choice of negative mass squared, which is above the Breitenlohner-Freedman bound for vectors in AdS spacetime, the asymptotic behavior of the fields as $z\to 0$ at the AdS boundary is
@@
The fact that the leading source term $\phi_{i}^{(0)}$ goes to zero as $z\to 0$ ensures the survival of the AdS asymptotics. [^1] As noted, the negative mass squared that we have chosen is above the Breitenlohner-Freedman bound for a massive vector field. This ensures that the scaling dimension $\Delta$ of the dual operator is real and the AdS vacuum is stable against classical perturbations. However, the scaling dimension corresponding to (6) is $\Delta=\frac{7}{4}$ which is below the CFT unitarity bound of $\Delta=2$ for a spin 1 operator in three dimensions. This means that our solutions do not describe deformations of a unitary CFT. It is unclear to us whether the boundary conditions (6), which include a source, are necessarily pathological from the bulk point of view. Certainly the equations of motion that we will be solving, following from the Ansatz (2) and (3), are well-behaved. The skeptical reader may note that we could also have chosen a relevant deformation with a positive mass squared and hence consistent with the unitarity bound. The qualitative behaviour of the solutions is not expected to be different. We have found the choice in the main text more convenient for illustrative numerical purposes — for instance, as we shall see shortly, Cauchy horizons are excluded a priori with a negative mass squared — even while perhaps the immediate holographic interpretation is less clear.
@@
We evolve the equations of motion in two steps. Firstly, we solve for the exterior geometry using variables $F$, $G$, $H$, and $\hat{\phi}_{i}\equiv\phi_{i}/z^{1/4}$. These variables admit a power series expansion in $\sqrt{z}$ and $(1-z)$ near the conformal boundary and black hole event horizon, respectively. We may then introduce a coordinate $\hat{z}\equiv\sqrt{z}$ with respect to which all functions admit analytic expansions near the conformal boundary. To solve the equation of motion, we employ a collocation method and discretise the system on a Gauss-Lobatto grid. Subsequently, the resulting algebraic equations are solved via a relaxation method using a standard Newton-Raphson routine. These methods are extensively reviewed in [53].
Solving for the interior geometry presents a significantly greater challenge. We first express the equations of motion in first-order form by introducing new variables $\tilde{G}\equiv G^{\prime}$ and $\tilde{\hat{\phi}}\equiv\hat{\phi}^{\prime}$. These variables are then evolved in the interior using the horizon values of all variables previously determined upon solving for the exterior geometry. To achieve long term stability, we use an adaptive implicit Runge-Kutta method of the Radau type [54]. Without the use of an adaptive method, these simulations would be nearly impossible to perform. Specifically, the required time step to accurately evolve all equations varies widely. The final challenge arises from the need to evolve the system over an extended period, involving extremely small numerical values. To address this issue, we employ octuple precision. This challenge proves to be the primary bottleneck in our computations, primarily due to its demanding memory requirements. A typical simulation, at its peak usage, can demand up to 150GB of RAM and over 10GB of storage.
@@ 3.3 The end of a Kasner epoch
$V_{n}=\text{max}\left(-h_{n},\frac{h_{n}-\sqrt{3}g_{n}}{2},\frac{h_{n}+\sqrt{3}g_{n}}{2}\right)\equiv\text{max}\left(V_{n}^{t},V_{n}^{x},V_{n}^{y}\right)\,.$ (23)
@@
These constraints follow from (12), (22) and (23). The following Fig. 4
Figure 4: The late time interior dynamics of $g(\rho)$ and $h(\rho)$ is confined to the triangle (24). The plots show three different illustrative trajectories, generated numerically using the full equations of motion (4) and (5). The straight lines correspond to Kasner epochs. An era is given by a sequence of epochs bouncing between two given walls. The earliest bounces extend slightly outside the triangle as the exponential walls have not yet become infinitely steep. The leftmost plot is zoomed in to a tip of the triangle, showing an era of many epochs. Left: $\phi_{t}^{(0)}\approx 1.49011\,T^{5/4}$, $\phi_{x}^{(0)}=3\phi_{t}^{(0)}$ and $\phi_{y}^{(0)}=4\phi_{t}^{(0)}$. Middle: $\phi_{t}^{(0)}\approx 2.88867\,T^{5/4}$, $\phi_{x}^{(0)}=2\phi^{(0)}_{t}$ and $\phi_{y}^{(0)}=4\phi^{(0)}_{t}$. Right: $\phi_{t}^{(0)}\approx 5.50918\,T^{5/4}$, $\phi_{x}^{(0)}=2\phi_{t}^{(0)}$ and $\phi_{y}^{(0)}=3\phi_{t}^{(0)}$.
shows three examples of the bounded evolution of the universe. It is clear from the figure that each Kasner epoch ends in a ‘bounce’ to a subsequent Kasner epoch. We will characterise such bounces in detail in the following sections.
While the potential (11) is not identical to the original Bianchi IX mixmaster potential from [1, 2, 3, 4], the walls in (24) are the same as the dominant walls in that case. In the Bianchi IX model the walls are curvature walls, whereas in our model — which has flat spatial slices and no curvature walls — they are electric walls. The sub-dominant curvature walls in the Bianchi IX dynamics, absent in our potential, are outside of the triangle (24) and are therefore irrelevant for the asymptotic interior dynamics. This is why we will recover the known recursion relations between Kasner epochs.
@@ 3.4 Reflection rule
$\left\{\begin{array}[]{c}p_{n}^{-}\to p_{n+1}^{0}\\ p_{n}^{0}\to p_{n+1}^{-}\\ p_{n}^{+}\to p_{n+1}^{+}\end{array}\right.\qquad\text{or}\qquad\left\{\begin{array}[]{c}p_{n}^{-}\to p_{n+1}^{+}\\ p_{n}^{0}\to p_{n+1}^{-}\\ p_{n}^{+}\to p_{n+1}^{0}\end{array}\right.\,.$ (31)
@@ 4.1 Era dynamics
$\xi_{N}\equiv\log\rho_{N}\qquad\text{and}\qquad\delta_{N}\equiv\frac{\Delta_{N}}{\rho_{N}}\,.$ (47)
@@ 5.1 The Wheeler-DeWitt equation
Here we have again considered late interior times so that the mass and cosmological constant terms in the action can be dropped. The momentum constraint is trivial in spatially homogeneous minisuperspace models.
@@
As commented in footnote 3 above, the minisuperspace WDW equation is only trustworthy to leading order in the semiclassical limit. Beyond that limit fluctuations of inhomogeneous modes contribute to the wavefunction and ordering ambiguities become relevant. Restricting, then, to highly excited states one has from (77) that
@@ 5.3 Congruence subgroups and automorphic forms
$x=0,\qquad x=1\qquad\text{and}\qquad\left(x-\frac{1}{2}\right)^{2}+y^{2}=\frac{1}{4}\,.$ (88)
@@ 6.1 Numerical method to determine large collections of eigenvalues
$\phi_{m}(\varepsilon_{n})=\left\{\begin{array}[]{cc}0&\text{if}\quad\sqrt{3}\pi m\leq\varepsilon_{n}\\ \sqrt{3\pi^{2}m^{2}-\varepsilon_{n}^{2}}-\varepsilon_{n}\arccos\left(\frac{\varepsilon_{n}}{\sqrt{3}\pi m}\right)&\text{if}\quad\sqrt{3}\pi m>\varepsilon_{n}\end{array}\right..$ (106b)
@@
As previously, the matrix $\mathbf{B}({\varepsilon}_{n})$, with components $B_{lm}({\varepsilon}_{n})$, is ill-conditioned. Finding suitable pre-conditioners might appear challenging, as the boundary condition (108) involves multiple Bessel functions that need to balance against each other. However, after some trial and error we have found that it is sufficient to define a matrix $\mathbf{B}^{\prime}({\varepsilon}_{n})$ just as in (106b), but with $m$ in (106b) replaced by $(m-1/2)$. For the Neumann case, computing the matrix $\mathbf{B}^{\prime}({\varepsilon}_{n})$ is more daunting, and we have only managed to accurately determine the first 1200 eigenvalues. Again, below, we will provide tests of our data suggesting that we have captured all the Neumann eigenvalues in the range $0<\varepsilon_{n}\lesssim 170.31(7)$.
@@ 6.2 Numerical scheme for finding $\varepsilon_{n}$ and few thousand $c^{n}_{m}$
With a given $\varepsilon_{n}$ known to some precision, we can change gears and use a different numerical method to determine $\varepsilon_{n}$ and the corresponding $c^{n}_{m}$ more accurately. In this method one solves for the eigenpair $\{\Psi_{n},\varepsilon_{n}\}$ directly on a discretised grid. First, we change to a new set of coordinates wherein $(x,y)$ are mapped to coordinates $(X,Y)$,
$X=x\quad\text{and}\quad Y=\frac{y}{\sqrt{1+y^{2}}}\,,$ (110)
@@
$X^{2}+\frac{Y^{2}}{1-Y^{2}}\geq 1\,,\quad 0\leq X\leq 1/2\,,\quad\text{and}\quad\sqrt{\frac{3}{7}}\leq Y\leq 1\,.$ (111)
@@ 6.3 Validating the numerical data
where the subscript D is for ‘Dirichlet’. The leading term in the expansion is determined by the area of the billiard and is independent of the boundary conditions or shape of the billiard. [^8] Locally one may consider the Laplace equation to be in flat space with eigenstates $e^{ik\cdot x}$ and eigenvalues $\mu=k^{2}$. Given a region of area $A$ one can, by the usual counting of plane wave states, fit $N=A\int d^{2}k/(2\pi)^{2}$ states into the region. Setting $d^{2}k=2\pi kdk$ and integrating up to $k=\sqrt{\mu}$ gives $N(\mu)=\frac{A}{4\pi}\mu$. The area of the small triangle defined by the interior of (96) is simply $\pi/6$.
@@
where the subscript N is to ‘Neumann’ and $\{A_{0},B_{0},C_{0}\}$ are to be fit. The first term in (114) is the same as for the Dirichlet sector. A simple $\chi^{2}$ fit yields
$A_{0}\approx 0.0010(3)\,,\quad B_{0}\approx-0.7043(6)\,,\quad\text{and}\quad C_{0}\approx 13.1661(2)\,.$ (115)
@@ 7.1 Density of states
where $A=\pi/6$ is the area of the small triangle defined in (96). We recalled the origin of this leading term in footnote 8 above. Although we shall work at fixed energy throughout, we can make a brief comment about the canonical ensemble. The density of states (117) corresponds to the entropy
@@ 7.4 Hecke operators
The matrices in (132) do not belong to $\Gamma(2)$, and hence it is not immediate that the action (131) maps the space of $\Gamma(2)$ automorphic forms to itself. We will now recall how this works.
@@
$\lambda_{p}^{n}d_{1}^{n}=d_{p}^{n}\quad(\text{Neumann})\quad\text{or}\quad\lambda_{p}^{n}d_{2}^{n}=d_{2p}^{n}\quad(\text{Dirichlet})\,.$ (141)
@@ 7.5 Waveform coefficients
$\mathrm{d}\mu(x)=\left\{\begin{array}[]{cc}\displaystyle\frac{1}{\pi}\sqrt{1-\frac{x^{2}}{4}}\,\mathrm{d}x&\quad\text{for}\quad|x|<2\\ 0&\text{otherwise}\end{array}\right.\,.$ (147)